OpenStudy (anonymous):

How do I multiply (5+x)(5+2x). Help is greatly needed

OpenStudy (lilg132):

5 * 5 = 25 5 * 2x = 10x x * 5 = 5x x * 2x = 2x^2 so 2x^2 + 5x + 10x + 25 which is now 2x^2 + 15x + 25 remember to click good answer

OpenStudy (anonymous):

Use the parrot beak technique the lines of the beak are the numbers you have to multiply: so 5*5 + x*5 + x*5 + x*2x